OpenStudy (inowalst):

\frac{ 8+2x }{ 4x }=\frac{ 3x }{ 4x }

OpenStudy (inowalst):

I did this:

OpenStudy (campbell_st):

ok... so start with a little factoring \[\frac{2(4 + x)}{4x} = \frac{3x}{4x}\] if you remove the common factors you get \[\frac{4 + x}{2x} = \frac{3}{4}\] now cross multiply \[4(4 + x) = 2x(3)\] now distribute and solve for x

OpenStudy (inowalst):

\[(4x)(8+2x)=(4x)(3x)\]\[32x+8x^2=12x^2\]\[32x+8x^2-8x^2=12x^2-8x^2\]\[32x-32x=4x^2-32x\]\[0=4x^2-32x\] I've done that so far.

OpenStudy (campbell_st):

here is my problem... looking at your work... x = 0 is a solution... which means you have on the left side \[\frac{8 + 2 \times 0}{4 \times 0} = \frac{3 \times 0}{4 \times 0}\] I think this is an issue

OpenStudy (inowalst):

\[4x=0\]\[\frac{ 4x }{ 4 }=\frac{ 0 }{ 4 }\]\[x=0\]

OpenStudy (campbell_st):

the problem I see is that dividing by zero is undefined...? when you use the original equation.

OpenStudy (inowalst):

\[\frac{ 8+2x }{ 4x }=\frac{ 3x }{ 4x }\]

OpenStudy (inowalst):

That's the original.

OpenStudy (leonardo0430):

;^)

OpenStudy (theraggedydoctor):

dot.